The resultant of two equal forces acting at a point at an angle 60 degrees is 6 N find the forces.

Grade:12th pass

2 Answers

Aksh
6 years ago
Let the two forces be F1 & F2, |F1|=|F2|=F|F1+F2|=√F×F+F×F+2Fcos60°=6N....(I) SQUARING BOTH SIDES F×F+F×F+2F(1/2)=362F×F+F =362F×F+F-36=0(F-4)(2F+9)=0F=4N or F=-9/2Discard -9/2 as it does not satisfy the equation (i)Therefore, the forces are of 4N each.
Aksh
6 years ago
Let the two forces be F1 & F2, |F1|=|F2|=F|F1+F2|=√F×F+F×F+2Fcos60°=6N....(I) SQUARING BOTH SIDES F×F+2F×F1/2=362F×F=36F=2√3The forces are of magnitude 2√3 N each